### v1

@@ A.3 SNAPSHOT Data Description
Sunrise
Time of sunrise (UTC)
Moon_phase
Moon phase value on a scale of 0-1(new moon-full moon)
Apparent_temp_max
Maximum apparent temperature of the day (°F)
Apparent_temp_min
Minimum apparent temperature of the day (°F)
Temperature_max
Maximum temperature of the day (°F)
Temperature_min
Minimum temperature of the day (°F)
Avg_cloud_cover
Percentage of sky covered by cloud on a scale of 0-1
Avg_dew_point
Average dew point temperature
Avg_humidity
Daily average value of humidity on a scale of 0-1
Avg_pressure
Average atmospheric pressure on the sea level (hPa)
Morning_pressure_change
Trinary value of pressure difference between midnight and noon (rising, falling, steady)
Evening_pressure_change
Trinary value of pressure difference between noon and midnight (rising, falling, steady)
Avg_visibility
Average visibility (meters)
weather_precip_probability
Precipitation probability
Temperature_rolling_mean
Rolling average of temperature
Temperature_rolling_std
Rolling standard deviation of temperature
apparentTemperature_rolling_mean
Rolling average of apparent temperature
apparentTemperature_rolling_std
Rolling standard deviation of apparent temperature
apparentTemperature_today_vs_avg_past
Difference between today’s apparent temperature and its rolling average
pressure_rolling_mean
Rolling average of pressure
pressure_rolling_std
Rolling standard deviation of pressure
apparentTemperature_today_vs_avg_past
Difference in today’s apparent temperature and its rolling average
pressure_rolling_mean
Rolling average of pressure
pressure_rolling_std
Rolling standard deviation of pressure
pressure_today_vs_avg_past
Difference between today’s pressure and its rolling average
cloudCover_rolling_mean
Rolling average of cloud cover
cloudCover_rolling_std
Rolling standard deviation of cloud cover
cloudCover_today_vs_avg_past
Difference between today’s cloud cover and its rolling average
humidity_rolling_mean
Rolling average humidity
humidity_rolling_std
Rolling standard deviation humidity
humidity_today_vs_avg_past
Difference between today’s humidity and its rolling average
@@
windSpeed_rolling_mean
Rolling average of wind speed
windSpeed_rolling_std
Rolling standard deviation of wind speed
windSpeed_today_vs_avg_past
Difference between today’s wind speed and its rolling average
precipProbability_rolling_mean
Rolling average of precipitation probability
precipProbability_rolling_std
Rolling standard deviation of precipitation probability
precipProbability_today_vs_avg_past
Difference between current precipitation probability and its rolling average
sunlight
Duration of sunlight (sec)
quality_of_day
Quality of the day defined in terms of 8 categories in the range -4, 4: clear=4, partly-cloudy=3, cloudy=2, wind=1, fog=-1, rain=-2, sleet=-3, snow=-4
avg_quality_of_day
Average value for quality_of_day
precipType
Type of precipitation as integer: None=0, Rain=1, Hail=2, Sleet=3, Snow=4, Other=5
max_precip_intensity
Maximum Precipitation volume (mm)
median_wind_speed
Median wind speed of the day (meter/sec)
median_wind_bearing
Median wind bearing of the day (degrees)
Appendix B
In this appendix, we provide more information on the experimental process from section 5. We provide model structures and parameters, and elaborate on the Reweighting method.
B.1 Model Structures
@@
B.2 Model Parameters/Implementation
We applied our bias-reducing multi-task learning method by developing MTL models to predict target and protected labels by which the baseline models are biased. We first trained the multi-task model for each of the combinations for all datasets. We assigned different loss weight ratios to the label pairs, prioritizing the prediction label over the protected label to ensure that the model prioritizes the prediction label and make certain that its uncertainty score is kept lower than the protected label prediction. For the ADULT dataset, the ratio was 4.5 to 0.25; for MIMIC-III, 5 to 0.5; and for SNAPSHOT, 6 to 0.25. We arrived at these combinations by experimenting with a few different combinations. For example, with the ADULT dataset, the uncertainty of the age label remained around 0.60 for varying ratios up to 4 to 1 and improved to 0.62 at ratios 4 to 0.25. For the SNAPSHOT dataset, we used Focal loss as the loss function for both predictions and assigned gamma values 2 and 0 for the chosen target and protected labels, respectively. We did this because setting gamma 0 reduces the relative loss for well-classified examples. We trained the MTL networks for 100 epochs for all the datasets and saved the weights at each epoch, making it a total of 100 weight combinations for each experiment run.
B.3 Reweighting
The Reweighting bias mitigation technique works by applying different weights to each group-label combination according to the conditional probability of label by the protected label.
@@ B.3 Reweighting
We assign to every tuple a weight according to its T and Label values. The balanced dataset is then created by sampling the original training data, replaced with the assigned weights. On this balanced dataset the dependency-free classifier is learned. The Reweighing technique can be seen as an instance of cost-sensitive learning Elkan (2001) in which, e.g., an object of label ‘+’ with T=t gets a higher weight, making an error for this object more "expensive". The pseudocode of the algorithm describing this Reweighing technique can be found in Calders et al. (2009) and Kamiran and Calders (2009).
Appendix C
In this appendix, we provide the saliency maps for the MIMIC-III and SNAPSHOT Experiments. Figures 15 to 23 correspond with the SNAPSHOT experiments. For these figures, the images on the left represent the saliency map for the initial baseline model for evening-sad-happy prediction, and those on the right are the maps for the model generated from our proposed method. Figure 24 was generated from the initial baseline model for IHM prediction for the MIMIC experiments, and Figure 25 for the model from our proposed method implementation with bias mitigation based on marital status.

### v2

@@ Appendix A. Data Description
Sunrise Time of sunrise (UTC)
Moon_phase Moon phase value on a scale of 0-1(new moon-full moon)
Apparent_temp_max Maximum apparent temperature of the day (°F)
Apparent_temp_min Minimum apparent temperature of the day (°F)
Temperature_max Maximum temperature of the day (°F)
Temperature_min Minimum temperature of the day (°F)
Avg_cloud_cover Percentage of sky covered by cloud on a scale of 0-1
Avg_dew_point Average dew point temperature
Avg_humidity Daily average value of humidity on a scale of 0-1
Avg_pressure Average atmospheric pressure on the sea level (hPa)
Morning_pressure_change Trinary value of pressure difference between midnight and noon (rising, falling, steady)
Evening_pressure_change Trinary value of pressure difference between noon and midnight (rising, falling, steady)
Avg_visibility Average visibility (meters)
weather_precip_probability Precipitation probability
Temperature_rolling_mean Rolling average of temperature
Temperature_rolling_std Rolling standard deviation of temperature
apparentTemperature_rolling_mean Rolling average of apparent temperature
apparentTemperature_rolling_std Rolling standard deviation of apparent temperature
apparentTemperature_today_vs_avg_past Difference between today’s apparent temperature and its rolling average
pressure_rolling_mean Rolling average of pressure
pressure_rolling_std Rolling standard deviation of pressure
apparentTemperature_today_vs_avg_past Difference in today’s apparent temperature and its rolling average
pressure_rolling_mean Rolling average of pressure
pressure_rolling_std Rolling standard deviation of pressure
pressure_today_vs_avg_past Difference between today’s pressure and its rolling average
cloudCover_rolling_mean Rolling average of cloud cover
cloudCover_rolling_std Rolling standard deviation of cloud cover
cloudCover_today_vs_avg_past Difference between today’s cloud cover and its rolling average
humidity_rolling_mean Rolling average humidity
humidity_rolling_std Rolling standard deviation humidity
humidity_today_vs_avg_past Difference between today’s humidity and its rolling average
@@
windSpeed_rolling_mean Rolling average of wind speed
windSpeed_rolling_std Rolling standard deviation of wind speed
windSpeed_today_vs_avg_past Difference between today’s wind speed and its rolling average
precipProbability_rolling_mean Rolling average of precipitation probability
precipProbability_rolling_std Rolling standard deviation of precipitation probability
precipProbability_today_vs_avg_past Difference between current precipitation probability and its rolling average
sunlight Duration of sunlight (sec)
quality_of_day Quality of the day defined in terms of 8 categories in the range -4, 4: clear=4, partly-cloudy=3, cloudy=2, wind=1, fog=-1, rain=-2, sleet=-3, snow=-4
avg_quality_of_day Average value for quality_of_day
precipType Type of precipitation as integer: None=0, Rain=1, Hail=2, Sleet=3, Snow=4, Other=5
max_precip_intensity Maximum Precipitation volume (mm)
median_wind_speed Median wind speed of the day (meter/sec)
median_wind_bearing Median wind bearing of the day (degrees)
[…]
Some participants in the SNAPSHOT dataset do not have race and ethnicity information, and we dropped samples for such participants when running experiments for race and ethnicity labels. In the end, we dropped 314 samples from the training data, reducing it from 4030 to 3716 samples, and the test samples were reduced from 1866 to 1715. We split the data into 70% training and 30% testing sets for all experiments.
Figure 10: SNAPSHOT Demographic Distribution
Appendix B.Experiments
In this appendix, we provide more information on the experimental process from Section 5. We provide details on baseline model implementations, model structures and parameters, and elaborate on the Reweighting method.
Baseline Model Implementations
For the ADULT dataset, we built a Convolutional Neural Network (CNN) with 2 dense layers and a dropout later with a dropout rate of 0.25 to predict binary income labels; $<=50$k as 0 and $>50$k as 1. We used a 3-fold cross-validation method and binary cross entropy as the loss function with an Adam optimizer. The structure of this model is shown below.
For the MIMIC-III dataset, we used a Long-Short Term Memory (LSTM)-based model proposed in Harutyunyan et al. [2019]. LSTM is a Recurrent Neural Network (RNN) which is a type of Artificial Neural Network (ANN) designed to capture long-term dependencies in sequential data (Harutyunyan et al. [2019]). To implement our model, we resampled the time series data into regularly spaced intervals. In the case of multiple measurements of the same variable in the same interval, we utilized the value of the last measurement. We used a 48-hour window for prediction, enabling the detection of patterns that may indicate changes in patient acuity, as proposed by Harutyunyan et al. [2019]. We selected the channel-wise LSTM model without deep supervision for our study based on its superior reported area under the receiver operating characteristic (AUROC) performance among the five developed non-MTL models by Harutyunyan et al. [2019]. The selected network consists of a bi-directional layer, an LSTM layer with 16 neurons, and a dropout layer with a dropout rate of 0.3 (structure shown in Appendix B). We used a batch size of 8, binary cross entropy as the loss function and Adam optimizer. We executed this for 100 epochs, reproducing the experiments from Harutyunyan et al. [2019].
The baseline model we used for predicting the labels for the SNAPSHOT dataset is a gated recurrent unit (GRU) model with 2 GRU layers and 2 dropout layers with a dropout rate of 0.5, 30 neurons and focal loss with a gamma value of 2, and an alpha value of 4. This model predicts 4 labels, namely: "morning happiness", "morning calmness", "evening happiness", and "evening calmness". We trained this model for 300 epochs with a batch size of 1000 and a focal loss function since there is a significant class imbalance in the target label for this dataset. Focal loss is an extension of the cross-entropy loss function that would down-weight easy examples and focus training on hard negatives (Lin et al. [2017]). It reduces the loss contribution from easy examples and increases the importance of correcting misclassified examples. The structure of this model is also shown in Appendix B. With 8 protected labels, which are gender, race, ethnicity, openness, conscientiousness, extraversion, agreeableness, and neuroticism, and 4 target labels, we conducted 32 sets of experiments to analyze the bias in this dataset.
Model Structures
@@
Proposed Method Implementation Details
We applied our bias-reducing MTL method by developing MTL models to predict target and protected labels by which the baseline models are biased. We first trained the MTL model for each of the combinations for all datasets. We assigned different loss weight ratios to the label pairs, prioritizing the prediction label over the protected label to ensure that the model prioritizes the prediction label and make certain that its uncertainty score is kept lower than the protected label prediction. For the ADULT dataset, the ratio was 4.5 to 0.25; for MIMIC-III, 5 to 0.5; and for SNAPSHOT, 6 to 0.25. We arrived at these combinations by experimenting with a few different combinations. For example, with the ADULT dataset, the uncertainty of the age label remained around 0.60 for varying ratios up to 4 to 1 and improved to 0.62 at ratios 4 to 0.25. For the SNAPSHOT dataset, we used Focal loss as the loss function for both predictions and assigned gamma values 2 and 0 for the chosen target and protected labels, respectively. We did this because setting gamma 0 reduces the relative loss for well-classified examples. We trained the MTL networks for 100 epochs for all the datasets and saved the weights at each epoch, making it a total of 100 weight combinations for each experiment run.
Data Manipulation on SNAPSHOT Dataset
To create a more significant imbalance for the gender label in the SNAPSHOT dataset, we manipulated the data in the following way:
Originally, the data contained samples from 181 male and 132 female participants. There was a total of 4030 train and 1866 test samples, and we removed the samples from 75 participants that identified as female from the training data, making it a total of 2660 samples for training. We removed participants with 14 or more samples in the train data (experiment S-1). We also experimented with removing samples from participants with 5 or more samples in the train data (experiment S-2) and those with 30 or more samples (experiment S-3), leaving 2349 and 3685 samples, respectively. The motivation is to induce an artificial data bias, otherwise known as a negative legacy, into our model by creating a higher imbalance in the representation of male and female participant samples in our training data. With this manipulation, we expect our model to not generalize well on the test data that still contains a comparable number of female samples as hypothesized by Gorrostieta et al. [2019], thereby inducing some bias in its predictions.
Reweighting
We compare our proposed method to a standard bias mitigation method called Reweighing introduced by Kamiran and Calders [2012]. Reweighing is a pre-processing bias mitigation technique. It involves applying appropriate weights to different tuples in the training dataset to make it discrimination free with respect to the protected label (Park et al. [2021]). It generates weights for the training examples in each (group, label) combination differently to ensure fairness before classification (Bellamy et al. [2018]). See the details of reweighting in Appendix B.
We apply the AI Fairness 360 implementation of this method to our processed data for all datasets and pass the weight vectors to our baseline model for all 3 datasets and calculated their fairness metric scores to compare to our proposed method.
The Reweighting bias mitigation technique works by applying different weights to each group-label combination according to the conditional probability of label by the protected label. It assigns different weights to various subgroups, determined by the relationship between the protected label and the outcome label. For example, if certain groups are underrepresented in positive outcomes, Reweighting increases their influence in the training process by assigning higher weights to these groups. This process aims to equalize the predictive performance across groups, mitigating any potential bias that the model may learn from imbalanced or skewed data.
@@
We assign to every tuple a weight according to its T and Label values. The balanced dataset is then created by sampling the original training data, replaced with the assigned weights. On this balanced dataset the dependency-free classifier is learned. The Reweighing technique can be seen as an instance of cost-sensitive learning (Elkan [2001]) in which, e.g., an object of label ‘+’ with T=t gets a higher weight, making an error for this object more "expensive". The pseudocode of the algorithm describing this Reweighing technique can be found in Calders et al. [2009] and Kamiran and Calders [2009].
Other Baseline Methods
We compare our proposed method’s performance on the ADULT and MIMIC-III datasets against the Fairness with Related Features (FairRF) method, introduced by Zhao et al. [2022]. The FairRF develops fair classifiers by leveraging features related to but not explicitly revealing sensitive attributes. The method focuses on minimizing the correlation between these related features and model predictions to ensure fairness. FairRF adjusts the importance weights of each feature and has shown effectiveness in enhancing classifier fairness while maintaining high accuracy across various real-world datasets.
Additionally, we review a fair algorithm called Adversarial Reweighted Learning (ARL) introduced by Lahoti et al. [2020] against our results for MIMIC-III. This approach focuses on identifying and mitigating biases that arise from non-protected features that are often correlated with sensitive attributes. The core idea of ARL is to use adversarial training techniques to re-weight the importance of features during the learning process. The adversarial component involves a dual-training setup where one model aims to predict the target outcome based on re-weighted features, while a second adversarial model attempts to predict the protected attribute from the same features. The goal is to adjust the weights of features in a way that the primary model maintains high accuracy on the target task, while the adversarial model fails to predict the protected attribute, thus ensuring that the predictions are fair and unbiased. This method was tested across various datasets and has shown to improve fairness metrics significantly without substantial loss in performance.
Appendix C. Saliency Maps for MIMIC-III and SNAPSHOT Datasets
In this appendix, we provide the saliency maps for the MIMIC-III and SNAPSHOT Experiments. Figures 14 to 22 correspond with the SNAPSHOT experiments. For these figures, the images on the left represent the saliency map for the initial baseline model for evening-sad-happy prediction, and those on the right are the maps for the model generated from our proposed method. Figure 23 was generated from the initial baseline model for IHM prediction for the MIMIC experiments, and Figure 24 for the model from our proposed method implementation with bias mitigation based on marital status.